Solve by method of separation of variable Î ”
1
Expert's answer
2021-10-12T15:40:04-0400

Solve by method of separation of variable

"ydx-xdy=0"

Solution:

"ydx-xdy=0
\\ \Rightarrow ydx=xdy
\\ \Rightarrow \dfrac {dx}x=\dfrac {dy}y"

On integrating both sides,

"\ln x+\ln C=\ln y
\\\Rightarrow \ln Cx=\ln y
\\\Rightarrow Cx=y"